OpenStudy (anonymous):

What Is The Value Of C-5/8C=20 C=?

OpenStudy (michele_laino):

please is your equation this? \[\frac{ c-5 }{ 8c }=20\]

OpenStudy (michele_laino):

if so, then multiply both sides of your equation by 8c, then write what you get, please

OpenStudy (anonymous):

Nope Not It

OpenStudy (michele_laino):

please, what is it then?

OpenStudy (michele_laino):

is this? \[C-\frac{ 5 }{ 8 }C=20\]

OpenStudy (anonymous):

Take Off The First C

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

so is this: \[-\frac{ 5 }{ 8 }C=20\]

OpenStudy (michele_laino):

if so, please multiply boith sides of your equation by 8

OpenStudy (anonymous):

so it will be Again -32

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (michele_laino):

@pootypie thank you very much!